# H-AMR FORGE’d in FIRE I: Magnetic state transitions, jet launching and radiative emission in super-Eddington, highly magnetized quasar disks formed from cosmological initial conditions

### v2

@@ 4.3 Caveats
  • The disk was initially evolved in the GIZMO code, which included different physics than the H-AMR code and different numerical treatment of some of the same physics. This is especially true for radiation. While both GIZMO and H-AMR use two-moment “M1” closure schemes (Levermore, 1984) for radiation, GIZMO uses a multiband method whereas H-AMR uses a grey (frequency-integrated) method. Additionally, the opacities in GIZMO are more sophisticated at low temperatures and have been well-tested in the interstellar medium in conditions more similar to the very outer regions of the disk and the photosphere.
@@
  • Our opacities assume that the radiation is thermal and in local thermodynamic equilibrium with the gas. In optically thin or marginally optically thick regions, these assumptions are inappropriate. While we still expect our effective temperatures at the photosphere to be correct at the order of magnitude level, they remain crude. This may also affect the high rate of conversion of the BZ jet’s Poynting flux into radiation, since this happens in the optically thin jet funnel. We also neglect dust opacities and partial ionization effects, which can both be important where the gas temperature is $<10^{4}\,{\rm K}$.
@@ 4.4 Summary
We report the results of two general-relativistic, radiation magnetohydrodynamics simulations of a quasar disk feeding a $1.3\times 10^{7}\,M_{\odot}$ SMBH, with $a=0$ and $0.9375$. We carried out the simulation using the H-AMR code. Our initial conditions were remapped from GIZMO, in which the disk was self-consistently formed within a galaxy (Hopkins et al., 2024b, c, a). This is the first time a quasar has been assembled from cosmological initial conditions and evolved down to the event horizon of the SMBH. We have specifically focused on the magnetic evolution of the inner accretion disk and the resulting radiative and outflow properties. We have also gleaned insights into the cosmological evolution of the SMBH. Our main findings are,
@@
Such strongly magnetized disks as we have studied here represent a frontier in accretion theory that is not well understood. The initial conditions we have used here represent a single data point, and a wider variety of studies – on both disk assembly from galaxy scales, as done here, and more controlled numerical experiments – are necessary to understand how generic this type of accretion flow is. There are several future directions that we can highlight. Firstly, the MHD turbulence mechanisms – especially in the outer disk – are not well understood. Yet, the resulting turbulence provides the reservoir of poloidal magnetic flux which dictates the magnetic state of the inner disk via advection physics that is also poorly understood. While we have not studied angular momentum transport and turbulence in this work, we plan to do so in the future. Secondly, it is also essential to understand the interplay of the MHD physics with thermodynamics, especially in the cold outer disk. The dominance of the toroidal field over the cold disk may be inhibiting the buoyant loss of toroidal flux (see also Squire et al., 2024). Also, the turbulence in the outer disk may be enhanced and/or strongly influenced by multi-phase gas physics and optically-thin cooling effects (Hopkins et al., 2024c) not modeled herein. Thirdly, related to the question of thermodynamics, it is necessary to study similar accretion flows in a wider range of Eddington ratios to understand how well our results connect to other types of AGN.
@@ B.2.2 Iterating through each block on the grid
The block boundary cleaning algorithm introduced in Section B.2.1 must be repeated for every block in the grid, but blocks share boundaries. If we first clean a block somewhere in our grid and then move to a neighbor, their shared boundary is already cleaned. So, as we “walk” through the grid, cleaning each block, there are less and less uncleaned boundaries remaining. We run the risk of some blocks having zero remaining boundaries to clean. Additionally, our grid has a non-trivial structure due to the combination of SMR and AMR. So, we must proceed carefully. We use the following approach,
@@ B.4 Scalar Divergence Cleaning
We also need to discretize the divergence of the magnetic field. Here and in the following equations, we write the discretized $\sqrt{-g}$ as $g_{i,j,k}$ for brevity, {widetext}
@@
$\frac{g_{i,j+1/2,k}B^{2}_{i,j+1/2,k}-g_{i,j-1/2,k}B^{2}_{i,j-1/2,k}}{\Delta x_{2}}+\frac{g_{i,j,k+1/2}B^{3}_{i,j,k+1/2}-g_{i,j,k-1/2}B^{3}_{i,j,k-1/2}}{\Delta x_{3}}\biggr{)}$
@@ B.5 Result of cleaning procedure
In Figure 16, we show the result of our cleaning process. In the top panels, we show a snapshot of $\vec{\nabla}\cdot\vec{B}\times\Delta x/|B|$ in the entire domain, before and after cleaning. Before cleaning, the normalized divergence reaches values $\mathcal{O}(10^{-1})$, which is expected given the truncation errors in the interpolation scheme. After cleaning, the normalized divergences reaches $\mathcal{O}(10^{-10})$ almost everywhere. Near the BH, where the field is essentially made up due to a lack of GIZMO data within $10\,r_{\rm g}$, the normalized divergence reaches $\mathcal{O}(10^{-7})$. We consider this an extremely good result. Additionally, we show the disk average (Eq. 7)) of the absolute value of the “physical” components of the magnetic field (Eq. 2) before and after interpolation. The difference is usually very small, with the largest discrepancies happening within $10\,r_{\rm g}$. This suggests that our interpolation of the GIZMO data is faithful.

### v4

@@ 4.3 Caveats
  • The disk was initially evolved in the GIZMO code, which included different physics than the H-AMR code and different numerical treatment of some of the same physics. This is especially true for radiation. While both GIZMO and H-AMR use two-moment “M1” closure schemes (Levermore, 1984) for radiation, GIZMO uses a multiband method whereas H-AMR uses a gray (frequency-integrated) method. Additionally, the opacities in GIZMO are more sophisticated at low temperatures and have been well-tested in the interstellar medium in conditions more similar to the very outer regions of the disk and the photosphere.
@@
  • Our opacities assume that the radiation is thermal and in local thermodynamic equilibrium with the gas. In optically thin or marginally optically thick regions, these assumptions are inappropriate. Additionally, since our radiation scheme is grey (frequency-integrated), we are forced to choose an averaging procedure for our opacities. This prevents our opacities from being accurate in all regimes; we have used an absorption average (McKinney et al., 2017), intended to be accurate in the highly optically thin or optically thick regimes, but this likely fall short in transitional regimes of moderate optical depth. While we still expect our effective temperatures at the photosphere to be correct at the order of magnitude level, they remain crude. This may also affect the high rate of conversion of the BZ jet’s Poynting flux into radiation, since this happens in the optically thin jet funnel. We also neglect dust opacities and partial ionization effects, which can both be important where the gas temperature is $<10^{4}\,{\rm K}$.
  • We have neglected radiation viscosity. This can be important in AGN disks since the radiation pressure far exceeds the gas pressure (e.g., Jiang et al., 2019a). However, in the super-Eddington disks we have simulated, radiation viscosity is less important because the mean free path of the photons is small (Jiang et al., 2019b). However, radiation viscosity may still play an important role in regions of moderate optical depth (i.e., where the mean free path is comparable to the local length scales of interest) such as the jet sheath or in lower-density, highly magnetized filaments, and deserves future study.
@@ 4.4 Summary
We report the results of two general-relativistic, radiation magnetohydrodynamics simulations of a quasar disk feeding a $1.3\times 10^{7}\,M_{\odot}$ SMBH, with $a=0$ and $0.9375$. We carried out the simulation using the H-AMR code. Our initial conditions were remapped from GIZMO, in which the disk was self-consistently formed within a galaxy (Hopkins et al., 2024c, d, a). This is the first time a quasar has been assembled from cosmological initial conditions and evolved down to the event horizon of the SMBH. We have specifically focused on the magnetic evolution of the inner accretion disk and the resulting radiative and outflow properties. We have also gleaned insights into the cosmological evolution of the SMBH. Our main findings are,
@@
Such strongly magnetized disks as we have studied here represent a frontier in accretion theory that is not well understood. The initial conditions we have used here represent a single data point, and a wider variety of studies – on both disk assembly from galaxy scales, as done here, and more controlled numerical experiments – are necessary to understand how generic this type of accretion flow is. There are several future directions that we can highlight. Firstly, the MHD turbulence mechanisms – especially in the outer disk – are not well understood. Yet, the resulting turbulence provides the reservoir of poloidal magnetic flux which dictates the magnetic state of the inner disk via advection physics that is also poorly understood. While we have not studied angular momentum transport and turbulence in this work, we plan to do so in the future. Secondly, it is also essential to understand the interplay of the MHD physics with thermodynamics, especially in the cold outer disk. The dominance of the toroidal field over the cold disk may be inhibiting the buoyant loss of toroidal flux (see also Squire et al., 2024). Also, the turbulence in the outer disk may be enhanced and/or strongly influenced by multi-phase gas physics and optically-thin cooling effects (Hopkins et al., 2024d) not modeled herein. Thirdly, related to the question of thermodynamics, it is necessary to study similar accretion flows in a wider range of Eddington ratios to understand how well our results connect to other types of AGN.
@@ B.2.2 Iterating through each block on the grid
The block boundary cleaning algorithm introduced in Section B.2.1 must be repeated for every block in the grid, but blocks share boundaries. If we first clean a block somewhere in our grid and then move to a neighbor, their shared boundary is already cleaned. So, as we “walk” through the grid, cleaning each block, there are fewer and fewer uncleaned boundaries remaining. We run the risk of some blocks having zero remaining boundaries to clean. Additionally, our grid has a non-trivial structure due to the combination of SMR and AMR. So, we must proceed carefully. We use the following approach,
@@ B.4 Scalar Divergence Cleaning
We also need to discretize the divergence of the magnetic field. Here and in the following equations, we write the discretized $\sqrt{-g}$ as $g_{i,j,k}$ for brevity,
@@
$\frac{g_{i,j+1/2,k}B^{2}_{i,j+1/2,k}-g_{i,j-1/2,k}B^{2}_{i,j-1/2,k}}{\Delta x_{2}}+$
$\frac{g_{i,j,k+1/2}B^{3}_{i,j,k+1/2}-g_{i,j,k-1/2}B^{3}_{i,j,k-1/2}}{\Delta x_{3}}\biggr{)}$
@@ B.5 Result of cleaning procedure
In Figure 16, we show the result of our cleaning process. In the top panels, we show a snapshot of $\vec{\nabla}\cdot\vec{B}\times\Delta x/|B|$ in the entire domain, before and after cleaning. Before cleaning, the normalized divergence reaches values $\mathcal{O}(10^{-1})$, which is expected given the truncation errors in the interpolation scheme. After cleaning, the normalized divergences reaches $\mathcal{O}(10^{-10})$ almost everywhere. Near the BH, where the field is essentially made up [^10] While this is not ideal, the near-BH region is small in volume, and the field added here is extremely sub-dominant to the reservoir of magnetic flux advected from large radii. This will wash away early transients. due to a lack of GIZMO data within $10\,r_{\rm g}$, the normalized divergence reaches $\mathcal{O}(10^{-7})$. We consider this an extremely good result. Additionally, we show the disk average (Eq. 7) of the absolute value of the “physical” components of the magnetic field (Eq. 2) before and after interpolation. The difference is usually very small, with the largest discrepancies happening within $10\,r_{\rm g}$. This suggests that our interpolation of the GIZMO data is faithful.